How many pounds of walnuts sell for $4.79 per pound must be mied with 4 pounds of cashews that sell for $4.90 per pound if the mixture is to sell for $4.84 per pound
Answer by checkley71(8403)

4.79x+4*4.90=(4+x)4.84
4.79x+19.6=19.36+4.84x
4.79x-4.84x=19.36-19.6
-.05x=-.24
x=-.24/.05
x=4.8 pounds of $4.79 walnuts are required.
proof:
4.79*4.8+19.6=(4+4.8)4.84
22.292+19.6=8.8*4.84
42.592=42.592